NvidiaIntely (nvidiaintely):

Challenge: Find the radius of the base of a cone with volume equal to 144π in3 and height equal to 12 in.

NvidiaIntely (nvidiaintely):

@Kevin

OpenStudy (kevin):

1/3 * pi * r * r = 144 * pi 1/3 * r * r = 144 r = 12 sqrt(3)

NvidiaIntely (nvidiaintely):

Thanks!